OpenStudy (anonymous):

What is 36x^2-84x+49 completely factored? I REALLY need help!!!

OpenStudy (ajprincess):

\(a^2x^2-2abx+b^2=(ax-b)^2\) \(36x^2-84x+49=6^2x^2-2*7*6x+7^2\) Do you see any similarities between these two @bambam423

OpenStudy (anonymous):

They kind of follow the same pattern. Sorta

OpenStudy (anonymous):

so that are your factors ] (6x-7)(6x-7)

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

its easier to ask yourself, looking at the last term, what times what is equal to 49. since 49 is a perfect square AND positive you know that its (?x-7) (?x-7) because a negative x a negative is a positive.

OpenStudy (anonymous):

Does that help a little?

OpenStudy (anonymous):

you can apply that to any equation you need factored. 4x^2+20x+25 what times what is 25? 5x5=25 (2x+5) (2x+5)

OpenStudy (anonymous):

Yes it does. I know some of it but our teacher is really bad and no one knows the answer! Thank you all! Now I understand better!

OpenStudy (anonymous):

Certainly :) im afraid i was a bit late to the party.